OpenStudy (anonymous):

What would be the most logical first step to solve this quadratic equation? (x+5)^2=36 a. Take the square root of both sides b. Subtract 36 from both sides so you can immediately factor c.Subtract five from both sides d. Make a difference of squares so you can immediately factor

OpenStudy (amistre64):

realizing that 6^2 = 36 so we have to subtract a 5 -5 +- 6

OpenStudy (anonymous):

i would use a) sensei?

OpenStudy (anonymous):

oooh this is a trick question thought up by some evil math teacher. i can just see it

OpenStudy (anonymous):

if you are a thinking human being you would write \[x+5=6\] or \[x+5=-6\] as a first step and then solve

OpenStudy (anonymous):

I would go with "a."

OpenStudy (anonymous):

i would call that "taking plus or minus the square root" but that is not what it says!

OpenStudy (anonymous):

i have to say that i don't understand the difference between b and d. anyone?

OpenStudy (anonymous):

a :) because a and are on opposite sides of the equation .

OpenStudy (anonymous):

A

OpenStudy (anonymous):

*b &c

OpenStudy (anonymous):

Thank you eveyone!

OpenStudy (anonymous):

@sensei realizing that \[6^2=36\] counts as a step!

OpenStudy (amistre64):

a is what it calls for ...

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

not without the sophistication that if \[x^2=36\] then \[x=6\] or \[x=-6\] and that is not "taking the square root" at least not literally. that is why i thought this might be a trick

OpenStudy (anonymous):

perhaps it is not a question from an evil math teacher. perhaps it is one to discuss, because there is really only one clear wrong answer. and even that isn't really clearly wrong if we don't count obviousness as a first step

OpenStudy (radar):

b and d is essentially saying the same thing. Looks like a is it.